# Łukasz Struski,  Jacek Tabor

### v1

@@ Abstract
Understanding the decisions made by image classification networks is a critical area of research in deep learning. This task is traditionally divided into two distinct approaches: post-hoc methods and intrinsic methods. Post-hoc methods, such as GradCam, aim to interpret the decisions of pre-trained models by identifying regions of the image where the network focuses its attention. However, these methods provide only a high-level overview, making it difficult to fully understand the network’s decision-making process. Conversely, intrinsic methods, like prototypical parts models, offer a more detailed understanding of network predictions but are constrained by specific architectures, training methods, and datasets.
In this paper, we introduce InfoDisent, a hybrid model that combines the advantages of both approaches. By utilizing an information bottleneck, InfoDisent disentangles the information in the final layer of a pre-trained deep network, enabling the breakdown of classification decisions into basic, understandable atomic components. Unlike standard prototypical parts approaches, InfoDisent can interpret the decisions of pre-trained classification networks and be used for making classification decisions, similar to intrinsic models. We validate the effectiveness of InfoDisent on benchmark datasets such as ImageNet, CUB-200-2011, Stanford Cars, and Stanford Dogs for both convolutional and transformer backbones.
Introduction and related works
Machine learning techniques, such as deep neural networks, have become essential tools in various applications including classification, image generation, speech recognition, and natural language processing. These techniques have achieved remarkable effectiveness, often rivaling human capabilities and sometimes surpassing them. A significant recent development in deep learning is the focus on model interpretability, giving rise to a subfield of artificial intelligence known as eXplainable AI (XAI) (Xu et al. 2019). Interpreting and understanding model learning is crucial in critical applications such as medicine, image recognition, autonomous driving, etc (Struski et al. 2024; Khan et al. 2001; Bojarski et al. 2017; Samek et al. 2021; Nauta et al. 2023b; Patrício, Neves, and Teixeira 2023).
Historically, the world of XAI is divided into two disjoint categories: post-hoc interpretability (Ribeiro, Singh, and Guestrin 2016; Lundberg and Lee 2017; Selvaraju et al. 2017), where we analyze the pre-trained model to explain its predictions, and inherently explained models (Chen et al. 2019; Böhle, Fritz, and Schiele 2022), where the aim lies in building networks which decisions are easy to interpret. Both of the above approaches have their advantages and disadvantages.
@@
Figure 1: Decision explanation constructed by InfoDisent for the pre-trained ViT (Dosovitskiy et al. 2020) feature space on the Agaric mushrooms image from the ImageNet dataset (left). In the prototype block (right) each row represents the prototypes (the corresponding channel number is on the right side of the block). The yellow boxes in each row show the activation of a given prototype, while in the first column, we show the activation of corresponding prototypes in the original image. Observe that we can informally trace the decision of ViT behind assigning the class Agaric to the image on the left to having a hat (channel 569), a white leg (channel 728), a reddish shine (channel 552), a strawberry texture (channel 297) and the appearance of ground with moss (channel 311).
@@
In the post-hoc methods, we interpret existing pre-trained network architectures. The commonly used methods like SHAP (Shapley 1951; Lundberg and Lee 2017), LIME (Ribeiro, Singh, and Guestrin 2016), LRP (Bach et al. 2015) or Grad-CAM (Selvaraju et al. 2017) provide in practice only feature importance which can be visualized as a saliency map that shows on which part of the image the model has focused its attention. This allows us to check if the model does not focus its attention outside of the object of interest (Ribeiro, Singh, and Guestrin 2016), however, it is in general not sufficient to really understand the reasons behind given predictions. Additionally, post hoc methods allow typically only local explanations (per the prediction of a given image), and do not allow to understand of the prerequisites to the given class.
@@
Inherently explained models
While post-hoc methods are easy to implement due to their non-intrusive nature, they often produce biased and unreliable explanations (Adebayo et al. 2018). To address this, recent research has increasingly focused on designing self-explainable models that make the decision process directly visible (Brendel and Bethge 2019; Alvarez Melis and Jaakkola 2018). Many of these interpretable solutions utilize attention mechanisms (Liu et al. 2021; Zheng et al. 2019) or exploit the activation space, such as with adversarial autoencoders (Guidotti et al. 2020). Among the most recent approaches, ProtoPNet (Chen et al. 2019) has significantly influenced the development of self-explainable models. It learns class-specific prototypes, similar to concepts, with a fixed number per class. The model classifies inputs by calculating responses from each class’s prototypes and summarizing these responses through a fully connected layer, providing explanations as a weighted sum of all prototypes. This method inspired the development of several other self-explainable models (Donnelly, Barnett, and Chen 2022; Rymarczyk et al. 2021, 2022; Wang et al. 2021; Nauta et al. 2023a; Donnelly, Barnett, and Chen 2022; Wang et al. 2021). Typically, in the prototypical parts model, the final decision of a given class is decomposed into the appearance of a few selected prototypes, which are similar to some strongly localized parts of some chosen images from the dataset. We see that in this approach the final class decision is split into a set of simpler and interpretable compounds.
Although intrinsic methods, like prototypical parts models, offer a more detailed understanding of network predictions as compared to post-hoc methods, they are constrained to specific architectures, training methods, and datasets [^1] In particular prototypical parts methods work well if the dataset consists of classes which are variations of objects of similar shape (like birds or cars), and consequently typically do not obtain a good prediction on diverse datasets like ImageNet..
@@
Thus in our opinion, one of the most important problems in XAI is whether it is possible to construct Post-hoc methods that would provide similar functionality to inherently interpretable ones. In particular, similarly to prototypical models, we would like to be able to split the final decisions into simpler interpretable components. Additionally, the ability to explain global predictions (what are the reasons for choosing a given class in general) is extremely desirable.
In this paper, we propose a hybrid model InfoDisent, whose aim is to take the best from both worlds and provide a partial positive answer to the above problem. The motivation of InfoDisent was to add the functionality of prototypical models to pre-trained architectures (we test InfoDisent both on convolutional and transformer architectures). To do so InfoDisent disentangles the information from the feature layer into separate prototype channels (similarly to the prototype-based model PIP-Net (Nauta et al. 2023a)). This is achieved by applying the information bottleneck, where we encourage the sparsity of the activation of a given prototype channel. Consequently, InfoDisent constructs interpretations of feature space of pre-trained models, which allow the decomposition of the final decision into simpler atomic components. Consequently, we can explain the decision behind a given class by looking at the most active prototype channels, see Figure 1.
Summarizing, the contributions of the paper are the following:
  • we propose InfoDisent, a hybrid model devoted to explainability, which lies in the border between Post-hoc and intrinsic models,
  • InfoDisent has similar applicability to prototypical parts models, but constructs interpretations for feature space of pre-trained networks,
  • the final class decision of InfoDisent can be traced back to easily interpretable prototypical channels.
InfoDisent: architecture
Our motivation comes from the general methodology of prototypical models, which can be described in the following way:
@@
Most interpretable models involve retraining certain parts of the CNN (Chen et al. 2019; Rymarczyk et al. 2022), whereas others, like PIP-Net (Nauta et al. 2023a), retrain the entire CNN. In contrast, our method utilizes a pre-trained CNN or transformer without further modification during training. We first employ a trainable orthogonal transformation $U$ on pixel space to enable the disentanglement of hidden features from feature maps [^2] To parametrize orthogonal maps, we restrict to those with positive determinants and use the formula $U=\exp(W-W^{T})$, where $\exp(\cdot)$ denote the matrix exponential (Hall and Hall 2013). We utilize the fact that the space of orthogonal matrices with positive determinants coincide with exponentials of skew-symmetric matrices (Shepard, Brozell, and Gidofalvi 2015). . To enforce the disentanglement we follow by an introduced sparse analogue of average pooling over channel $K$ given by
@@
Observe that to compute $\mathrm{mx\_pool}(K)$ we need to know only the highest positive and negative pixel values of $K$, contrary to $\mathrm{avg\_pool}$, where all the pixel values are needed. Subsequently, we identify sparse representations (superpixels) within the channels that contribute positively or negatively to predictions. This enables us to generate heatmaps akin to Grad-CAM (Selvaraju et al. 2017) without necessitating a backward model step, as shown in Figure 5(a). Importantly, unlike Grad-CAM, our technique supports the visualization of negative heatmaps, resembling the LRP (Bach et al. 2015) method that requires a backward pass in a neural network. Our method operates solely during the forward step.
@@
The crucial consequence of InfoDisent that is disentangles the channels making them interpretable. Thus, similarly to PiPNet (Nauta et al. 2023a), we identify channels as prototypes. To illustrate a given prototype channel, we present five images from the training dataset on which the activation of the channel is the greatest, see Figure 3, where we present consecutive prototypes for the pre-trained ResNet-50 model. This follows from the fact that for better interpretability of model decisions, it is beneficial for humans to be presented from 4 to 9 concepts (Rymarczyk et al. 2022). Formally, similarly as in prototypical models, as the prototypical part, we understand the part of the image corresponding to pixels in feature space with maximal activation, marked by the yellow box in  Figure 3. Observe that the presented prototypes seem consistent with each other, and could be well interpreted.
Figure 3: The image shows prototypes from channels 689 to 694 in a trained ResNet-50 model on the ImageNet dataset. Each row displays the 5 most significant prototypes from a single channel. The prototypes are highlighted in yellow boxes.
@@
Figure 4: Sample examples illustrating how the ResNet-50 model, trained on the ImageNet dataset, explains the input images using prototypes.
@@
Observe that, unlike Grad-CAM, our heatmaps also illustrate negative activations, akin to the Layer-Wise Relevance Propagation (LRP) method. While LRP can effectively highlight both positive and negative contributions to the model’s decision, it often faces challenges such as complexity in implementation and sensitivity to model architecture variations. Our method addresses these issues by providing clear and interpretable visualizations. Sample results using this explanation method are presented in Figure 5(b).
@@
Experiments and Results
This section outlines the experiments conducted to compare our approach with current state-of-the-art methods, highlighting the most significant results. Our experiments utilized a variety of datasets, including well-established benchmarks for evaluating the explainability of artificial models: The Caltech-UCSD Birds-200-2011 (CUB-200-2011) (Wah et al. 2011), Stanford Cars (Krause et al. 2013), and Stanford Dogs (Khosla et al. 2011), as well as the newly introduced FunnyBirds (Hesse, Schaub-Meyer, and Roth 2023). Additionally, we incorporated the full ImageNet (Russakovsky et al. 2015) dataset for the first time in the class of prototypical networks.
Additional results and details concerning these experiments are provided in the Supplementary Materials.
@@
Appendix C Additional results
@@
InfoDisent employs prototypes, similar to the approach used in PiPNet (Nauta et al. 2023a), to explain individual decisions made on an image. Below, we describe the process for identifying prototypes for a given input image. An average pooling operation makes the aggregation of information from individual channels. The outcome of this operation is a scalar for each channel $K$, computed as
@@
(as explained in the main paper). This dense representation, formed by aggregating all the channels, is then processed by a linear layer that outputs logits, see LABEL:fig:appendix_architecture. The logits can be represented in a format similar to the output of a convolutional layer, as illustrated by $V_{+}$ and $V_{-}$ in LABEL:fig:appendix_architecture. This approach maintains the pictorial structure of the logits, allowing us to extract individual channels.
@@
Complementing our experiments with widely used CNN models, we investigated the performance of transformer architectures, specifically VisionTransformer (ViT-B/16) (Dosovitskiy et al. 2020), SwinTransformer (Swin-S) (Liu et al. 2022a). Their results are visualized in Figures 19 and 20. Our analysis reveals that transformer models concentrate on smaller image regions than CNNs. Nevertheless, both model types generate interpretable prototypes that offer insights into the input image content.
@@
Figures 24 and 25 present example heatmaps generated by our model, resembling those produced by Grad-CAM (Selvaraju et al. 2017). Our method, rooted in representational channels, simplifies heatmap generation by accumulating activations from all prototypes (positive and negative) across all channels. Leveraging the information bottleneck principle, our approach yields more focused heatmaps compared to traditional methods.

### v2

# Łukasz Struski 1,  Dawid Rymarczyk 1,2,  Jacek Tabor 1

@@ Abstract
In this work, we introduce InfoDisent, a hybrid approach to explainability based on the information bottleneck principle. InfoDisent enables the disentanglement of information in the final layer of any pretrained model into atomic concepts, which can be interpreted as prototypical parts. This approach merges the flexibility of post-hoc methods with the concept-level modeling capabilities of self-explainable neural networks, such as ProtoPNets. We demonstrate the effectiveness of InfoDisent through computational experiments and user studies across various datasets using modern backbones such as ViTs and convolutional networks. Notably, InfoDisent generalizes the prototypical parts approach to novel domains (ImageNet).
Introduction
Deep neural networks have demonstrated performance that matches or even surpasses human capabilities across various domains, such as image classification and generation, speech recognition, and natural language processing. Despite their impressive achievements, these networks often operate as ”black boxes”, offering little insight into the reasoning behind their decisions (rudin2019stop).
This lack of transparency poses significant challenges, particularly in high-stakes applications such as medicine and autonomous driving, where understanding a model’s decision-making process is critical (bojarski2017explaining; khan2001classification; nauta2023anecdotal; patricio2023explainable; samek2021explaining; struski2024multiple). To address this issue, the subfield of artificial intelligence known as eXplainable AI (XAI) (xu2019explainable) has emerged, focusing on making AI systems more interpretable and trustworthy.
XAI methods can be classified into two major categories: post-hoc methods and inherently interpretable (ante-hoc) methods (rudin2019stop). Post-hoc methods, such as GradCAM (selvaraju2017grad), are highly flexible because they can be applied to any neural network architecture. However, they are often unreliable (adebayo2018sanity), provide local explanations, and fail to reveal the characteristics of specific classes. In contrast, ante-hoc methods, such as ProtoPNet (chen2019looks), offer concept-level explanations by identifying prototypical parts from the training dataset. While this approach enhances interpretability, it is
@@ Introduction
Figure 1: Decision explanation constructed by InfoDisent for the pre-trained ViT feature space on the Agaric mushrooms image from the ImageNet. We can trace the decision of ViT behind assigning the class Agaric to the image on the left to having a hat (569), a white leg (728), a reddish shine (552), a strawberry texture (297) and the appearance of ground with moss (311). Note that, the prototype block (right) each row represents the prototypical part (the corresponding channel number). The yellow boxes in each row show the activation of a given prototypical part, while in the first column, we show the activation of corresponding prototypical parts in the original image.
limited to fine-grained classification tasks and specific architectures, such as CNNs in the case of ProtoPNet. Consequently, these methods do not fully leverage the potential of large pretrained models, such as Vision Transformers (ViTs) (dosovitskiy2020image).
To address this, we propose InfoDisent, a novel XAI approach that leverages information disentanglement in the final layer of any pretrained model. In InfoDisent, each channel encapsulates a single atomic concept, interpretable as prototypical parts – similar to the approach used in PIPNet (nauta2023pipnet). This is achieved through the application of an information bottleneck, in which we enforce activation sparsity in each prototypical channel. As a result, InfoDisent provides both local and global-level explanations in the form of human-interpretable concepts, as illustrated in Figure 1. Furthermore, InfoDisent offers significant flexibility, as it can be seamlessly applied to any pretrained architecture, including ViTs and convolutional networks.
We demonstrate the effectiveness of InfoDisent through both computational experiments and user studies. Our method is benchmarked on five datasets, including ImageNet, a challenging dataset where prototypical parts-based approaches have not been generalized to. Additionally, results from the user study indicate that InfoDisent provides a competitive level of explanation understanding compared to other prototypical parts-based methods, while offering the advantage of flexibility in its application to any model backbone. We make the code available.
Our contributions can be summarized as follows:
  • We propose InfoDisent, a hybrid XAI model that combines the interpretability of ante-hoc methods with the flexibility of post-hoc approaches.
  • To the best of our knowledge, InfoDisent is the first model to provide prototypical, part-like interpretations for the feature space of any pretrained network.
  • We validate the effectiveness of InfoDisent in terms of both accuracy and user understanding through extensive experimental evaluations.
Related Works
Research in XAI can be divided into two disjoint categories: post-hoc interpretability (lundberg2017unified; ribeiro2016should; selvaraju2017grad), where we analyze the pre-trained model to explain its predictions, and inherently explained models (bohle2022b; chen2019looks), where the aim lies in building networks which decisions are easy to interpret. Both of the above approaches have their advantages and disadvantages which we discuss in the following paragraphs.
@@
In the post-hoc methods, we interpret existing pre-trained network architectures. The commonly used methods such as SHAP (lundberg2017unified; shapley1951notes), LIME (ribeiro2016should), LRP (bach2015pixel) and Grad-CAM (selvaraju2017grad) provide in practice only feature importance which can be visualized as a saliency map that shows on which part of the image the model has focused its attention. This allows us to check if the model does not focus its attention outside of the object of interest (ribeiro2016should), however, it is in general not sufficient to really understand the reasons behind given predictions. Additionally, post-hoc methods allow typically only local explanations (per the prediction of a given image), and do not allow to understand of the prerequisites to the given class.
Inherently explained models
While post-hoc methods are easy to implement due to their non-intrusive nature, they often produce biased and unreliable explanations (NEURIPS2018_294a8ed2). To address this, recent research has increasingly focused on designing self-explainable models that make the decision process directly visible (brendel2018approximating; NEURIPS2018_3e9f0fc9). Many of these interpretable solutions utilize attention mechanisms (liu2021visual; zheng2019looking) or exploit the activation space, such as with adversarial autoencoders (Guidotti_Monreale_Matwin_Pedreschi_2020). Among the most recent approaches, ProtoPNet (chen2019looks) has significantly influenced the development of self-explainable models. It learns class-specific prototypes, similar to concepts, with a fixed number per class. The model classifies inputs by calculating responses from each class’s prototypes and summarizing these responses through a fully connected layer, providing explanations as a weighted sum of all prototypes. This method inspired the development of several other self-explainable models (donnelly2022deformable; nauta2023pipnet; pach2024lucidppn; rymarczyk2021protopshare; rymarczyk2022interpretable; rymarczyk2023icicle; wang2021interpretable). Typically, in the prototypical parts-based model, the final decision of a given class is decomposed into the appearance of a few selected prototypes, which are similar to some strongly localized parts of some chosen images from the training dataset. We see that in this approach the final class decision is split into a set of simpler and interpretable concepts.
@@
Our motivation comes from the general methodology of prototypical models, which can be described as follows:
@@
Most interpretable models involve retraining certain parts of the CNN (chen2019looks; rymarczyk2022interpretable), whereas others, like PIP-Net (nauta2023pipnet), retrain the entire CNN. In contrast, our method utilizes a pre-trained CNN or transformer without further modification during training. We first employ a trainable orthogonal transformation $U$ on pixel space to enable the disentanglement of hidden features from feature maps [^1] To parametrize orthogonal maps, we restrict to those with positive determinants and use the formula $U=\exp(W-W^{T})$, where $\exp(\cdot)$ denote the matrix exponential (hall2013lie). We utilize the fact that the space of orthogonal matrices with positive determinants coincide with exponentials of skew-symmetric matrices (shepard2015representation). . To enforce the disentanglement we follow by an introduced sparse analogue of average pooling over channel $K$ given by
@@
Observe that to compute $\mathrm{mx\_pool}(K)$ we need to know only the highest positive and negative pixel values of $K$, contrary to $\mathrm{avg\_pool}$, where all the pixel values are needed. Subsequently, we identify sparse representations (superpixels) within the channels that contribute positively or negatively to predictions. This enables us to generate heatmaps akin to Grad-CAM (selvaraju2017grad) without necessitating a backward model step, as shown in Figure 5(a). Importantly, unlike Grad-CAM, our technique supports the visualization of negative heatmaps, resembling the LRP (bach2015pixel) method that requires a backward pass in a neural network. Our method operates solely during the forward step.
@@
The crucial consequence of InfoDisent that is disentangles the channels making them interpretable. Thus, similarly to PiPNet (nauta2023pipnet), we identify channels as prototypes. To illustrate a given prototype channel, we present five images from the training dataset on which the activation of the channel is the greatest, see Figure 3, where we present consecutive prototypes for the pre-trained ResNet-50 model. This follows from the fact that for better interpretability of model decisions, it is beneficial for humans to be presented from 4 to 9 concepts (rymarczyk2022interpretable). Formally, similarly as in prototypical models, as the prototypical part, we understand the part of the image corresponding to pixels in feature space with maximal activation, marked by the yellow box in  Figure 3. Observe that the presented prototypes seem consistent with each other, and could be well interpreted.
Figure 3: The image shows prototypes from channels 689 to 692 in a trained ResNet-50 on the ImageNet. Each row displays the 5 most significant patches from a single prototypical channel. The prototype’s activations are highlighted by yellow boxes.
@@
Figure 4: Exemplary explanation (hen) for ResNet-50 backbone provided by InfoDisent in a form of prototypical parts.
@@
Observe that, unlike Grad-CAM, our heatmaps also illustrate negative activations, akin to the Layer-Wise Relevance Propagation (LRP) method. While LRP can effectively highlight both positive and negative contributions to the model’s decision, it often faces challenges such as complexity in implementation and sensitivity to model architecture variations. Our method addresses these issues by providing clear and interpretable visualizations. Sample results using this explanation method are presented in Figure 24.
@@
Numerical Experiments
This section outlines the experiments conducted to compare our approach with current state-of-the-art methods, highlighting the most significant results. Our experiments utilized a variety of datasets, including well-established benchmarks for evaluating the explainability of artificial models: The Caltech-UCSD Birds-200-2011 (CUB-200-2011) (wah2011caltech), Stanford Cars (krause20133d), and Stanford Dogs (khosla2011novel). Additionally, we incorporated the full ImageNet (russakovsky2015imagenet) dataset for the first time in the class of prototypical networks.
Additional results, including ablation study and analysis of inter-channel correlation, and details concerning those experiments are provided in the Supplementary Materials.
Table 2: Classification accuracy on full CUB-200-2011, and Stanford Dogs datasets by competing approaches using different CNN backbones. For each dataset and backbone, we boldface the best result in the class of interpretable models.
@@
Appendix D Additional results
@@
InfoDisent employs prototypes, similar to the approach used in PiPNet [nauta2023pipnet], to explain individual decisions made on an image. Below, we describe the process for identifying prototypes for a given input image. An average pooling operation makes the aggregation of information from individual channels. The outcome of this operation is a scalar for each channel $K$, computed as
@@
(as explained in the main paper). This dense representation, formed by aggregating all the channels, is then processed by a linear layer that outputs logits, see Figure 2. The logits can be represented in a format similar to the output of a convolutional layer, as illustrated by $V_{+}$ and $V_{-}$ in Figure 2. This approach maintains the pictorial structure of the logits, allowing us to extract individual channels.
@@
Complementing our experiments with widely used CNN models, we investigated the performance of transformer architectures, specifically VisionTransformer (ViT-B/16) [dosovitskiy2020image], SwinTransformer (Swin-S) [liu2022swin]. Their results are visualized in Figures 19 and 20. Our analysis reveals that transformer models concentrate on smaller image regions than CNNs. Nevertheless, both model types generate interpretable prototypes that offer insights into the input image content.
@@
Figures 24 and 25 present example heatmaps generated by our model, resembling those produced by Grad-CAM [selvaraju2017grad]. Our method, rooted in representational channels, simplifies heatmap generation by accumulating activations from all prototypes (positive and negative) across all channels. Leveraging the information bottleneck principle, our approach yields more focused heatmaps compared to traditional methods.
More details on user study
Each worker was paid €2.00 for completing a short 20-question survey. The survey questions were randomly composed, so the specific questions differed between participants. The participants were gender-balanced and ranged in age from 18 to 60. They were given 30 minutes to complete the survey.
To ensure data quality, we excluded responses where users selected the same answer for all questions. Surveys were repeated until we obtained 60 valid responses. Figure 26 and Figure 27 illustrate example questions used in both user studies.
Before starting the survey, participants were provided with an example and detailed instructions to familiarize them with the study setup, including the explanation composition and visualization. The distribution of answers is summarized in Tables 6 and Table 7.
Table 6: Distribution of answers in user study on user confidence in model’s prediction based on explanation.
Dataset Fairly … correct Somewhat … correct Somewhat … incorrect Fairly … incorrect
CUB 585 316 208 91
ImageNet 449 248 189 314
Table 7: Distribution of answers in user study on prototypical part disambiguity.
Dataset Correct Incorrect
CUB 776 424
ImageNet 712 488
Is InfoDisent reducing channel correlation?
To assess whether our model reduces channel correlation, we computed the RV coefficient, a standard measure of linear correlation between vectors. In 14 experiments (across 3 datasets and 7 backbones), our model outperformed the baseline in 9 cases (smaller RV coefficients), performed worse in 2 cases, and showed equivalent results in 3 cases as presented in Table 8. Despite not having a loss term explicitly aimed at reducing the RV coefficient, our model generally demonstrates either smaller or equal RV coefficients compared to the baseline.
Table 8: RV correlation between channels for InfoDisent and baseline models.
Dataset Model InfoDisent Baseline
Cars ConvNeXt Tiny 6.7 6.9
DenseNet121 6.2 6.9
ResNet34 6.0 7.1
ResNet50 6.2 7.6
CUB-200-2011 ConvNeXt Tiny 5.5 5.5
DenseNet121 7.2 7.9
ResNet34 6.5 7.2
ResNet50 4.9 7.8
ImageNet ConvNeXt Large 8.7 3.0
DenseNet121 8.4 4.5
ResNet34 5.2 6.0
ResNet50 3.0 6.5
SwinV2-S 2.5 2.5
ViT-B-16 9.9 9.9
@@
Figure 26: An exemplary question from the user study on user confidence.
Figure 27: An exemplary question from the user study on disambiguity of prototypical parts.